write the equation of the line that goes through the point (0,3/2) and is perpendicular to the line 5y=3x-4. the equation has to be in standard form. i dont know how to get rid of the fraction in the points.
Answer by Alan3354(69443)

write the equation of the line that goes through the point (0,3/2) and is perpendicular to the line 5y=3x-4. the equation has to be in standard form
---------
Find the slope, m1, of 5y=3x-4
5y=3x-4
y = (3/5)x - 4/5
m1 = 3/5
--------
Lines perpendicular will have a slope that's the negative inverse, -5/3
m2 = -5/3
y = mx + b thru (0,3/2)
3/2 = (-5/3)*0 + b
b = 3/2
y = (-5/3)x + 3/2
6y = -10x + 9
10x + 6y = 9